OpenStudy (anonymous):

Find the exact value of the expression. sec [ sin^-1 (-(√3)/2)]

hartnn (hartnn):

lets find the value of sin^-1 (-(√3)/2) first, let \(\sin^{-1} (-(√3)/2) = y \) so, \(-\sqrt3/2=\sin y\) do you know for which angle, value of sin y is -(√3)/2 ?

OpenStudy (anonymous):

5π/3 or 4π/3? o:

hartnn (hartnn):

you can use any value, you'll get the same answer in the end ^_^

hartnn (hartnn):

so, sec [..that angle...] =... ?

OpenStudy (anonymous):

Uhh, I'm not sure o;

hartnn (hartnn):

ok, its y=5pi/3 so, sec[5pi/3] =... ?

OpenStudy (anonymous):

1/2?

hartnn (hartnn):

1/2 or 2 ?

OpenStudy (anonymous):

Should I be putting it in the calculator? Or looking at the unit circle o;

hartnn (hartnn):

if you can use calculator, then its great...just plug in that

OpenStudy (anonymous):

It keeps saying error o: Let me try a different one

OpenStudy (anonymous):

I can't on the test o; Why does it keep saying domain error?

hartnn (hartnn):

what did you plug in calculator exactly ?

OpenStudy (anonymous):

cos^-1 (5π/3)

hartnn (hartnn):

why! ? sec [ sin^-1 (-(√3)/2)] we let y=sin^-1 (-(√3)/2) so, we need to find sec y we got y= 5pi/3 so, we just need sec (5pi/3) did you get this ?

OpenStudy (anonymous):

Yess, cos^-1 is the same as sec o;

hartnn (hartnn):

not at all!

OpenStudy (anonymous):

cos^-1 is not sec...if you want sec you have to put 1/[cos(x)]

hartnn (hartnn):

\(\cos x = 1/\sec x\) so, you try \(1/\cos[5\pi/3]\)

OpenStudy (anonymous):

And then I get 1.0042

OpenStudy (anonymous):

Change the mode...you have it in degree mode right now. It needs to be in radian

OpenStudy (anonymous):

Oo okay! Thank you guys very much (:

hartnn (hartnn):

welcome ^_^